please solve ASAP. :(


Find the equation of the plane which passes through the line of intersection of the planes x+y−2z = 1 and 2x+y−4z = 3 and which is perpendicular to the plane x+y+z = 1


1
Expert's answer
2021-06-30T09:18:40-0400

"x+y-2z-1=0\ \ and\ \ 2x+y-4z-3=0"


Equation of plane passing through the line of intersection of the above two planes is

"(x+y-2z-1)+\lambda(2x+y-4z-3)=0\\\therefore x(1+2\lambda)+y(1+\lambda)+z(-2-4\lambda)-1-3\lambda=0"


This plane is perpendicular to "x+y+z-1=0" "\implies" Normals are also perpendicular

Direction ratios of normals of new plane "a_1,b_1,c_1" are "(1+2\lambda),(1+\lambda),(-2-4\lambda)" respectively

Direction ratios of normals of plane x+y+z-1=0 "a_2,b_2,c_2" "are\ 1,1,1" respectively


Since the planes are perpendicular,

So, "a_1a_2+b_1b_2+c_1c_2=0"

"\therefore 1+2\lambda+1+\lambda-2-4\lambda=0"

"\therefore -\lambda=" 0 "\implies \lambda=0"


Substituting "\lambda=0\" in equation of line , we get

"\implies x+y-2z-1=0"

This is the required equation of line.